OpenStudy (anonymous):

P(A or B) = .4 and P (A and B) +.1, Find P(A) + P(B)

OpenStudy (anonymous):

P(A or B) = P(A) + P(B) - P(A and B) therefore 4 = P(A) + P(B) - 1 P(A) + P(B)=3

OpenStudy (anonymous):

P(A) + P(B)=4+1=5

OpenStudy (anonymous):

o...sry you are right...

OpenStudy (anonymous):

They both look right!? Idk!!

OpenStudy (anonymous):

If they are mutually exclusive then Salmon answer is right, if they are not then you are right. But it does not say on my homework assignment.

OpenStudy (kc_kennylau):

@salmon @eliassaab do you notice the little dot in ".4" and ".1"

OpenStudy (raden):

not sure P(A) + P(B) = 3 or more we knowed that maximum value of P(A) = 1, also for P(B)

OpenStudy (anonymous):

So close! I think I am just going to use one of their answers. I am not sure, thanks for all help!

OpenStudy (anonymous):

P(A or B) = P(A) + P(B) - P(A and B) therefore 0.4 = P(A) + P(B) - 0.1 P(A) + P(B)=0.4+0.1=0.5 This should be the most appropriate answer satisfying all the above comments

OpenStudy (anonymous):

Thank you!!